In the Bible, Exodus Chapter 18 describes how Moses was overwhelmed as the single judge of Israel because of the large number of trivial cases that were brought before him. His father-in-law, Jethro, recommended a hierarchical system of appellate courts where the lowest-level judge had responsibility for 10 citizens. Five judges of 10 sent the difficult cases that they could not resolve to a judge of 50 citizens. Two judges of 50 were under a judge of 100, and 10 judges of 100 were under a judge of 1,000. The judges of 1,000 citizens reported to Moses, who had to decide only the most difficult cases. If a population were exactly 2,000 citizens (excluding judges), what would be the total population, including Moses, all the judges, and citizens?
Answer #1

=> 131 Judges are required to handle 1000 citizens and most difficult cases will be reported to Moses by the Judge who handles 1000 citizens.

Since there are 2000 citizens, so we need one more judge who handles 1000 citizens in the same hierarchy.

=> 131+131 = 262 Judges are required to handle 2000 citizens.

=> the total population, including Moses, all the judges, and citizens = Citizens + Judges + Moses

= 2000 + 262 + 1 = 2263 Ans.

=> Go through this and let me know if you have any doubt in this!

=> Please do the positive rating if you are satisfied with our response ;)

Thank you ;)